OpenStudy (winsteria):

SOMEONE PLEASE HELP ME IF YOU DO I'LL GIVE YOU MEDAL AND FAN please help!!! What value of x makes the equation true? –(9 + (–17)) = –9 + x x =???

OpenStudy (anonymous):

x equal 17

OpenStudy (stefrheart):

lol first before i say anything is that 1d?

OpenStudy (stefrheart):

-9+-17=-9+x -26=-9+x x=-17

OpenStudy (anonymous):

- ( 9 + (-17 )) = - (9-17 ) = -9 +17 therefore x = 17

OpenStudy (stefrheart):

lol its just a copy of itself!

OpenStudy (stefrheart):

oh yeah forgot about that other parenthesis! sorry! its 17

OpenStudy (winsteria):

okay goodness thank both of you so much and yes it is 1D lol :)

OpenStudy (stefrheart):

hahaha welcome even though i made a mistake! And I was a huge directioner but kinda lost my touch!

OpenStudy (winsteria):

Oh really that's kinda sad omg what made you loose your touch ? :(

OpenStudy (stefrheart):

Here i explained it to you on my message